OpenStudy (anonymous):

use the sum of cubes to factor. be sure to factor out any common factors x^3-64

OpenStudy (anonymous):

notice the 4 is a \root \therefore \[x^3-64=(x-4)Q(x)\] Using long division we get \[x^3-64=(x-4)(x^2+4x+16)\]